(1 point) Use the Laplace transform to solve the following initial value problem: y6y9y 0,with y(0) 1, y (0) = -4 First, usin Please answer the blamnks. Thank you.
Answer #1

a y+(-6)y +gyzo y,0)=1 ,yllo)=ry y-byrgyzo 1910)=1 ,y(o)=-4 Take Laplace on both side Ily) -69 141)+2294)=0. Now a ly)6-5= A (5-3) + B for rood 873 3-5 =A-O+B > 11=-2] = 3-5= AB-3A+B thon Take inverse Yaplace we get (800) = 316--214-) te at ye